OpenStudy (anonymous):

Solve the equation lg(5x + 10) + 2 lg 3 = lg(4x + 12).

OpenStudy (anonymous):

use log a+log b =log ab

OpenStudy (anonymous):

lg(5x + 10) + 2 lg 3 = lg(4x + 12). log (5x + 10) *3^2=log (4x + 12) [ if loga = log b then a=b ] 9(5x+10)=4x+12 45x+90 = 4x+12 or 45x-4x=12 -90 or 41x =-78 hence x=-78/41

OpenStudy (anonymous):

Thanks! :)

OpenStudy (anonymous):

welcome